OpenStudy (anonymous):

[6.03] What is the value of the y variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −6 6 −3 3 −6

OpenStudy (shamim):

just multiply ur 1st equation by 2 and multiply ur 2nd equation by 3 then subtract those 2 equations

OpenStudy (anonymous):

what

OpenStudy (calculusxy):

Why don't you do the Guess, Check, and Revise method? What you just do is just plug in the integers towards the variables of x and y.

OpenStudy (shamim):

u hv 2 equations r \[3x+2y=6....................(1)\]

OpenStudy (shamim):

\[2x+3y=-6.........................(2)\]

OpenStudy (calculusxy):

I think that the answer to the variable is (-6).

OpenStudy (shamim):

if u multiply ur equation (1) by 2 u will get\[6x+4y=12...................(3)\]

OpenStudy (anonymous):

so the answer 3??

OpenStudy (shamim):

anyway do the math first

OpenStudy (calculusxy):

Oh never mind my answer.

OpenStudy (shamim):

then u hv to multiply ur equation (2) by 3, then u will get\[6x+9y=-18......................(4)\]

OpenStudy (anonymous):

I suck at math.

OpenStudy (shamim):

now subtract (4) frm equation (3)

OpenStudy (calculusxy):

Don't think it like that. Math seems difficult at first, but when you have the confidence, you can understand the math.

OpenStudy (anonymous):

1?

OpenStudy (shamim):

y=-6

OpenStudy (calculusxy):

Yes that's what I had just said. I figured that out by filling in the -6 in the x value and also the y value.

OpenStudy (calculusxy):

Anyhow, if there was two answers to the x function then it wouldn't be made into a function. This equation would have been a quadratic equation, when you have two answers.

OpenStudy (anonymous):

oh sorry i been working on math all night i feel so dum :(